OpenStudy (anonymous):

Find all solutions of the equation. sin x = -(radical3)/2

OpenStudy (anonymous):

\[\sin x = -\frac{\sqrt 3}{2}\]

OpenStudy (jdoe0001):

\(\bf sin(x)=-\cfrac{\sqrt{3}}{2}\implies sin^{-1}[sin(x)]=sin^{-1}\left(-\cfrac{\sqrt{3}}{2}\right)\\ \quad \\ \measuredangle x==sin^{-1}\left(-\cfrac{\sqrt{3}}{2}\right)\)